the sum of the ages of a and b is 45 years.after 10 years a age is twise that of b age.find a age?
Found 2 solutions by Susmitha, stanbon:
Answer by Susmitha(21)

given sum of the ages of a and b is 45
so, a+b=45
then given after ten years a age is twice that of b
so, a=2b
2b+b=45
3b=45
b=45/3
b=15
-----
a+ 15=45
a=45-15
a=30
--------

Answer by stanbon(75887)

the sum of the ages of a and b is 45 years.after 10 years a age is twise that of b age.find a age?
--------------------------
Equations:
a + b = 45
a+10 = 2(b+10)
---
Substitute for a and solve for "b":
45-b + 10 = 2b+20
3b = 35
b = 11 2/5 (age of b)
------
Solve for "a":
a = 45-b
a = 33 3/5 (age of a)
---------
Cheers,
Stan H.
===================